BellSouth Telephone Company wants to know how many telephone numbers are possible for each code. If the digits 1 and 0 cannot be used in the first two positions, how many seven-digits phone numbers are possible for each code?
Answer by stanbon(75887)

BellSouth Telephone Company wants to know how many telephone numbers are possible for each code. If the digits 1 and 0 cannot be used in the first two positions, how many seven-digits phone numbers are possible for each code?
---------
Ans: 9^2*10^5 = 81*100000 = 8,100,000
=============
Cheers,
Stan H.
-------------